5y=7x+22
x=-6y+17

Answer by mananth(16946)

5y=7x+22
x=-6y+17
Re write the equations


-7 x + 5 y = 22 .............1
1 x + 6 y = 17 .............2
Eliminate y
multiply (1)by -6
Multiply (2) by 5
42 x -30 y = -132
5 x + 30 y = 85
Add the two equations
47 x = -47
/ 47
x = -1
plug value of x in (1)
-7 x + 5 y = 22
7 + 5 y = 22
5 y = 22 -7
5 y = 15
y = 3